3rd RESCEU Symposium ‘Particle Cosmology’ Program

10 th. Nov. (Mon)

17:00– Registration 18:00–20:00 Welcome Party

11 th. Nov. (Tue)

9:10–9:15 Opening Address Sato, K.(Chairman of Organizing Committee) 9:15–10:45 (Chair. Yoshimura, M.)

Kofman, L. Reheating and Preheating after Inflation (40+5)
Boyanovksky, D  Nonperturbative Dynamics in a New Inflation Model
and Cosmological Perturbations  (40+5)

10:45–11:05 Coffee Break 11:05–12:10 (Chair. Kodama, H.)

Yoshimura. M.  Elementary Processes in Cosmology  (40+5)
Kim, S. P.  Aspects of Quantum Inflationary Cosmology  (15+5)

12:10–13:40 Lunch 13:40–15:55(Chair. Yokoyama, J.)

Tanaka, T.  One-Bubble Open Inflation Scenario and Quantum Field in an Open Universe (40+5)
Kawasaki, M.  New Inflation in Supergravity  (40+5)
Singh, A.  Instabilities in Axion Fluctuations  (15+5)
Kasuya, S.  Topological Defect Formation during Preheating  (20+5)

15:55–16:25 Coffee Break 16:25–18:10 (Chair. Sasaki, M.)

Vachaspati, T. Topological Defects: Fossils from the Early Universe (40+5)
Yokoyama, J.  On the Formation of Nontopological Strings
at the Electroweak Phase Transition  (15+5)
Fodor, Z.  Electroweak Phase Transition on Four-Dimensional Lattices (15+5)
Kim, Y.  Black Cosmic String  (15+5)

12 th. Nov. (Wed)

9:10–10:55 (Chair. Sato, K.)

Olive, K.  The Current status of BBN: Theory and Observations (40+5)
Levshakov, S. Primordial D Abundance Measurements (15+5)
Kohri, K.  Primordial Nucleosynthesis and Neutrino Cosmology (15+5)
Mathews, G. J.  Comment on Nonstandard Primordial Nucleosynthesis (15+5)

10:55–11:15 Coffee Break 11:15–12:20 (Chair. Yanagida, T.)

Murayama, H. Supersymmetric Cosmology (40+5)
Oshimo, N.  Electroweak Baryogenesis and Neutron Electric Dipole Moment
from Supersymmetry  (15+5)

12:20–14:00 Lunch 14:00–15:35(Chair. Hosoya, A.)

Frieman, J. Inflation and Large-scale Structure (40+5)
Piotrkowska, K.  Renormalisation Group and criticality in Einstein’s Universe (25+5)
Morikawa, M.  Renormalization Group Method in Newtonian Cosmology (15+5)

15:35–16:05 Coffee Break 16:05–17:35 (Chair. Kawasaki, M.)

Olinto, A.  Cosmic Magnetic Fields and Ultra-High Energy Cosmic Rays (40+5)
Kajita, T. Results from Super-Kamiokande – first 1 year – (40+5)

18:00–20:00 Banquet

13 th. Nov. (Thu)

9:10–10:30 (Chair. Tomimatsu, A.)

Ichinose, S.  Weyl Anomaly in 2D Dilaton Gravity and Hawking Radiation (15+5)
Hotta, K.  The Information Loss Problem of Black Hole
and the First Order Phase Transition in String Theory (15+5)
Koga, J.  Transformation of generalized black hole thermodynamics
into the Einstein frame  (15+5)
Torii, T. Black Hole Thermodynamics and Stability (15+5)

10:30–10:50 Coffee Break 10:50–12:10(Chair. Maeda, K.)

Hawking, S. W.  The Evaporation of Primordial Black Holes  (50)
Bousso, R.  Evaporation of Cosmological Black Holes  (20+10)

12:10–13:40 Lunch 13:40–15:05 13:40-15:05(Chair. Morikawa, M.)

Barrow, J. D.  Varying G and Other Constants  (40+5)
Ng, K.-W.  Cosmology from Broken Scale Symmetry  (15+5)
Spaans, M.  Topology and Homotopy in Grand Unified Theories (15+5)

15:05–15:25 Coffee Break 15:25–16:20 (Chair. Yanagida, T.)

Schramm, D.N.  Particle-Cosmology into the 21st Century  (50+5)